Question 6 of 12 (1 point) Use the average bond energies in the text to estimate the enthalpy changes of the following reactions: N2(g)+O2(g) → 2 NO Number k.J/mol C(s)+CO2(g) → 2 CO(g) Note: The heat of sublimation of graphite, C(s), is 719 kJ/mol . Number kJ/ mol
Answer #1

Energy is required to break the bonds and energy is released when the new bonds are formed . The net sum of these bond energies is called as enthalpy change of the reaction .

For the first reaction :

The bonds to be broken are :

N triple bond N ( B.E. = 941 KJ/mol)

O=O ( B.E. = 495 KJ/mol)

New bonds formed =

two N= O ( B.E. = 607 KJ/mol)

Delta H = 941 + 495 + (2 x -607)

= 222 KJ/mol

For the second reaction :

Bonds broken are :

two C=O = 799 KJ/mol

Bonds formed are :

two CO bonds = 1077 KJ / mol

And since the heat of sublimation is +719 KJ /mol

the delta H = 719 + (2 x 799) + (2 x -1077)

= 163 KJ/mol